OpenStudy (anonymous):

If the equation y = 6(1.6)^x is plotted on a semilog scale graph, then the linearized equation will be log(y) = m x + b where m = ? b = ?

OpenStudy (anonymous):

m=1.6, b=log6

OpenStudy (anonymous):

log base 6? or log(6)?:)

OpenStudy (anonymous):

log(6)

OpenStudy (anonymous):

thanks!:D can u show steps please:D! i can prob figure it out later though

OpenStudy (anonymous):

y = 6(1.6)^x this is what u have taking log to both sides log y=log6+log(1.6)^x =log 6 +x log 1.6